for a game you pay 2$ to play and roll 1 die. If the number is a 1 or 2 you win 9$. If the number is a 6, you win 3$. If the number is anything else, you receive nothing. FInd the expected value (the fair price to charge to play).
Answer #1

P($(9 - 2)) = P($7) = P(1 or 2) = 2/6 = 1/3

P($(3 - 2)) = P($1) = P(6) = 1/6

P(- $2) = P(3, 4, 5) = 3/6 = 1/2

Hence,

Expected value = $7 (1/3) + $1 (1/6) - $2 (1/2) = $3/2 = $1.5